For the equation of the line both in point-slope and slope-intercept form going through the point (3,-2) and parallel to the line with the equation -2y=6x-3
Answer by ankor@dixie-net.com(22740)

For the equation of the line both in point-slope and slope-intercept form going through the point (3,-2) and parallel to the line with the equation -2y=6x-3
:
Put the equation into the slope intercept form to obtain the slope
-2y = 6x - 3
divide both sides by -2
y = -3x + 1.5
slope is -3, parallel lines have the same slope
Find the equation of the line in the point slope form where
m = -3
x1 = 3
y1 = -2
y - (-2) = -3(x - 3)
y + 2 = -3(x - 3); is the point/slope form
y = -3x + 9 - 2
y = -3x + 7; is the slope/intercept form